### v1

@@ 7.3. Navigating Epistemic Risks in Operationalizing Positionality
However, translating complex human positionality into computational prompts introduces consequential epistemic risks. As highlighted by the inherent brevity of user profiles, there is a tension between making positionality computationally tractable and preserving the nuance of a researcher’s whole lived experience. As we move toward systems that model user perspectives, there is a danger of algorithmic essentialism, which reduces a researcher to a set of flat characteristics or stereotypes. Large language models, which rely on probabilistic correlations, may inadvertently generate reductive or essentializing attributions (e.g., suggesting a researcher’s interpretation is solely defined by a single demographic trait rather than their holistic expertise). This further risks reifying stereotypes or entrenching assumptions of how people with certain identities or perspectives interpret data, undermining the researcher’s authority and contextual information.
@@
Operationalizing and surfacing user perspectives in this way enables systems to facilitate more subtle interactions in high-stakes tasks, such as academic peer review. Despite showing signs of enhancing peer review quality (Thakkar et al., 2025), current research primarily focuses on efficiency (Chen et al., 2025; Hossain et al., 2024). But a system that models the reviewer’s perspective could address the challenge of interpretive diversity. For a meta-reviewer, such a system could contextualize conflicting assessments by highlighting how reviewers’ differing theoretical lenses shaped their critiques. However, our findings suggest a critical constraint for such future tools: to avoid reductive profiling, they must remain tentative and non-prescriptive, surfacing potential associations between diverse backgrounds, and act only to stimulate human deliberation.
@@ 8. Limitations
Our study and system have several limitations that bound our claims. First, the evaluation involved 12 experienced qualitative researchers working in pairs within a single, time-limited session on a synthetic dataset. This setting lets us observe rich, situated use of Reflexis and provides interaction-level insights into reflexivity features, but it does not capture the non-linear, longitudinal dynamics of real projects that unfold over months, nor the influence of institutional constraints, shifting research questions, or evolving team relationships. As such, our findings speak to initial use and perceived affordances rather than long-term change in practice.
@@
Finally, our operationalization of positionality is necessarily partial. Reflexis represents positionality as brief, self-authored profiles that are summarized into prompts. This makes positionality computationally tractable but risks flattening complex, intersectional, and evolving positions into relatively stable descriptors. We did not systematically examine potential harms or discomfort this might create for marginalized researchers or in sensitive domains, so we interpret our results as evidence of both promise and risk rather than as validation of positionality-aware prompting as a general best practice.
@@ 9. Future Work
Several directions follow from this work. First, we plan to conduct longitudinal deployments of Reflexis with teams using the system throughout the lifecycle of research projects. Such studies can examine how reflexive streams, provenance views, and positionality-focused collaboration are adapted, domesticated, or abandoned over time, and whether they ultimately reshape how findings are constructed and reported.
@@ 10. Conclusion
This paper introduces Reflexis, a collaborative workspace that deeply embeds researcher reflexivity, code evolution, and principled disagreement into the analytical workflow. Through a paired-analyst study with 12 qualitative researchers, we showed how integrating in-situ reflexive streams, analytical provenance, and disagreement-centered collaboration can encourage a more granular, deliberate, and methodologically transparent practice than is supported by many current toolsets and workflows. Our findings exemplify how AI can be used to scaffold reflexive and dialogic practices. At the same time, the study surfaces important risks and limits of operationalizing positionality in tools, such as prompts that resonate can anchor productive dialogue, while generic or misaligned framings risk flattening complex identities and underplay reflexive work. We offer Reflexis as a design probe towards the challenges of positionality in collaborative qualitative analysis. By intentionally designing for deliberation rather than efficiency, our work contributes a principled approach to building human-AI systems for interpretive knowledge work. The insights from Reflexis can guide the future tools for thought that honor and augment, rather than automate, the deep, personal processes of human sensemaking and knowledge-seeking.
@@ Acknowledgements.
We acknowledge and thank the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2024-04348 and RGPIN-2024-06005]. This work was also supported by the National Science Foundation under Award No. 2402428. Furthermore, we extend a heartfelt thanks to our study participants, without whom this work would not have been possible.
@@ 6 P11 MSc Student HCI 2 3
Table 2. User Study Participant Demographic. Participants rated their comfort level with RTA on a 7-point Likert scale, ranging from 1 (Extremely uncomfortable) to 7 (Extremely comfortable).
Group # ID RTA Experience (simplfied) View on Topic of Analysis (simplified)

### v2

@@ 7.3. Navigating Epistemic Risks in Operationalizing Positionality
However, translating complex human positionality into computational prompts introduces consequential epistemic risks. As highlighted by the inherent brevity of user profiles, there is a tension between making positionality computationally tractable and preserving the nuance of a researcher’s entire lived experience. As we move toward systems that model user perspectives, there is a danger of algorithmic essentialism, which reduces a researcher to a set of flat characteristics or stereotypes. Large language models, which rely on probabilistic correlations, may inadvertently generate reductive or essentializing attributions (e.g., suggesting a researcher’s interpretation is solely defined by a single demographic trait rather than their holistic expertise). This further risks reifying stereotypes or entrenching assumptions of how people with certain identities or perspectives interpret data, undermining the researcher’s authority and contextual information.
@@
Operationalizing and surfacing user perspectives in this way enables systems to facilitate more subtle interactions in high-stakes tasks, such as academic peer review. Despite showing signs of enhancing peer review quality (Thakkar et al., 2025), current research primarily focuses on efficiency (Chen et al., 2025; Hossain et al., 2024). But a system that models the reviewer’s perspective could address the challenge of interpretive diversity. For a meta-reviewer, such a system could contextualize conflicting assessments by highlighting how reviewers’ differing theoretical lenses shaped their critiques. However, our findings suggest a critical constraint for such future tools: to avoid reductive profiling, they must remain tentative and non-prescriptive, surfacing potential associations between diverse backgrounds, acting only to stimulate human deliberation.
@@ 8. Limitations
Our study and system have several limitations that bound our claims. First, the evaluation involved 12 experienced qualitative researchers working in pairs within a single, time-limited session on a synthetic dataset. This setting let us observe rich, situated use of Reflexis and provided interaction-level insights into reflexivity features, but it did not capture the non-linear, longitudinal dynamics of real projects that unfold over months, nor the influence of institutional constraints, shifting research questions, or evolving team relationships. As such, our findings speak to initial use and perceived affordances rather than long-term change in practice.
@@
Finally, our operationalization of positionality is necessarily partial. Reflexis operationalizes positionality as brief, self-authored profiles that are summarized into prompts. This makes positionality computationally tractable but risks flattening complex, intersectional, and evolving positions into relatively stable descriptors. We did not systematically examine potential harms or discomfort this might create for marginalized researchers or in sensitive domains, so we interpret our results as evidence of both promise and risk rather than as validation of positionality-aware prompting as a general best practice.
@@ 9. Future Work
Several directions follow from this work. First, we plan to conduct longitudinal deployments of Reflexis with teams using the system throughout the lifecycle of research projects. Such studies can examine how reflexive streams, provenance views, and positionality-focused collaboration are adapted, incorporated, or abandoned over time, and whether they ultimately reshape how findings are constructed and reported.
@@ 10. Conclusion
This paper introduces Reflexis, a collaborative workspace that deeply embeds researcher reflexivity, code evolution, and principled disagreement into the analytical workflow. Through a paired-analyst study with 12 qualitative researchers, we showed how integrating in-situ reflexive streams, analytical provenance, and disagreement-centered collaboration can encourage a more granular, deliberate, and methodologically transparent practice than is supported by many current toolsets and workflows. Our findings exemplify how AI can scaffold reflexive and dialogic practices. At the same time, the study surfaces important risks and limits of operationalizing positionality in tools, while effective prompts can anchor productive dialogue, generic or misaligned framings risk flattening complex identities and underplaying reflexive work. We offer Reflexis as a design probe towards the challenges of positionality in collaborative qualitative analysis. By intentionally designing for deliberation rather than efficiency, our work contributes a principled approach to building human-AI systems for interpretive knowledge work. The insights from Reflexis can guide the future tools for thought that honor and augment, rather than automate, the deep, personal processes of human sensemaking and knowledge-seeking.
@@ Acknowledgements.
We extend a heartfelt thanks to our study participants, without whom this work would not have been possible. We acknowledge and thank the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2024-04348 and RGPIN-2024-06005]. This work was also supported by the National Science Foundation under Award No. 2402428.
@@ 6 P11 MSc Student HCI 2 3
Table 2. User Study Participant Demographics. Participants rated their comfort level with RTA on a 7-point Likert scale, ranging from 1 (Extremely uncomfortable) to 7 (Extremely comfortable).
Group # ID RTA Experience (simplified) View on Topic of Analysis (simplified)